If Jennifer deposits $11,000 into a long-term Certificates of Deposit account that pays an annual interest rate of 2.7%, compounded continuously. Find the amount in the account after 8 years. Also, find how long it takes for her account to reach $20,000.
Answer by Boreal(15235)

P=Poe^rt for continuous compounding, where r=0.027 and t=8. Note, the rule of 72 would say 72/2.7 is the doubling time, which is about 27 years.
=11000*e^(.027*8)=10000*e^.216=$13652.13 This is as expected not even close to doubling.
20000=11000e(.027*t)
1.818=e^(.027t)
ln both sides, without rounding, then divide by 0.027
0.5978=0.027*t
t=22.14 years